OpenStudy (anonymous):

Imagine working with a younger sibling or student who is having trouble multiplying by 9s. Create an algorithm to help the student learn how to do this easily. Then notate whether the given algorithm is equivalent or not. You may need to use outside resources to help you.

OpenStudy (shaik0124):

9*1 = 10 - 1 = 9 9*2 = 20 - 2 = 18 9*3 = 30 - 3 = 27 9*4 = 40 - 4 = 36 9*5 = 50 - 5 = 45 9*6 = 60 - 6 = 54 9*7 = 70 - 7 = 63 9*8 = 80 - 8 = 72 9*9 = 90 - 9 = 81 9*10 = 100 -10 = 90 9*11 = 110 -11 = 99 9*12 = 120 -12= 108 9*13 = 130 - 13 = 117 9*14 = 140 - 14 = 126 The formula for 9 * N or (9N): Add a zero to N and minus N -or- multiply N by 10 and minus N 10N - N 9*200 = 2000 - 200 = 1800 9*999 = 9990 - 999 = 8991

OpenStudy (shaik0124):

got it?

OpenStudy (shaik0124):

is there anything more i can help u with?